OpenStudy (texmechanica):

Consider the system of equations: 3x+4y=67 and 5x-y=35 Edmundo used the multiplication property of equality so that the y-terms would be eliminated when the equations were added. What was the resulting equation?

OpenStudy (texmechanica):

@imqwerty Please help. I will fan and medal

OpenStudy (texmechanica):

Oh, I left something out.

OpenStudy (retireed):

3x+4y=67 multiplication property of equality so that the y-terms would be eliminated 5x-y=35 <----- Multiply by 4 20x - 4y = 140 3x+ 4y = 67 Add 23x = 207

OpenStudy (retireed):

Looks complete to me. Unless you want to solve for x or y.